OpenStudy (anonymous):

Please walk me through this problem! I understand that you sub in the 2.5 but how do you deal with all the decimals?

OpenStudy (lgbasallote):

you'll be posting the problem right? o.O

OpenStudy (anonymous):

Consider that the age, x, of a unicorn in human equivalent years can be given by the formula f(x) = - 0.001518x4 + 0.097326x3 – 1.2367x2 + 10.46x + 2.914. When a unicorn is 2.5 years old, what is its age in human equivalent years? What about when it is 12 years old?

OpenStudy (anonymous):

\[ f(x) = - 0.001518x^4 + 0.097326x^3 – 1.2367x^2 + 10.46x + 2.914\] \[ f(2.5) = - 0.001518(2.5)^4 + 0.097326(2.5)^3 – 1.2367(2.5)^2 + 10.46(2.5) + 2.914\] Now you just plug into the calculator, and you get your answer. If you are wondering about what: \[ (2.5)^4\] for instance means, it meas: 2.5*2.5*2.5*2.5= 39.0625 similarly: \[ (2.5)^3\] means: 2.5*2.5*2.5= 15.62500 and so on...

OpenStudy (anonymous):

that's where I was having trouble understanding. I was trying to multiple then add on the exponent Thanks!

OpenStudy (anonymous):

f(x)=23.132, do I divide off the f(2.5)?